OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (anonymous):

2x+10=5y 2x/5+2=y y=mx+c m is your slope m=2/5

OpenStudy (anonymous):

-5y=-2x-10 (-5y/-5)=(-2x/-5)-(10/-5) y=2x/5 + 2 2/5 is the slope.

OpenStudy (raden):

anyway, u just know what formula to find slope of line ax+by=c, m=-a/b

OpenStudy (calculusfunctions):

@michaelr35, how do find the slope? There are two approaches: i). you could rewrite the linear equation in slope-intercept form (y = mx + b), where the slope is m and the y-intercept is b. OR ii). If the linear equation is written in implicit form (x and y are both on one side of the equation), then the slope m is the negative of the coefficient of x divided by the coefficient of y. Try it and let me know if you require further assistance.